OpenStudy (anonymous):

Can someone actually work this problem out step by step with me? I know the answer already (i got it wrong on a quiz) and I just want to know how to do it step by step. The sum of two consecutive integers is at least 46. What is the least possible pair of integers?

OpenStudy (texaschic101):

do you know what consecutive means ?

OpenStudy (anonymous):

Umm kind of. Im doing all the quiz corrections right now, this test was from a long time ago, so i kind of forget.. But i can look it up!

OpenStudy (texaschic101):

consecutive means one right after the other....such as 1 and 2, or 4 and 5. so lets let 1 number be (x) and the other number be (x + 1). The sum of 2 consecutive numbers is at least 46.. x + (x + 1) >= 46 --- combine like terms 2x + 1 >= 46 -- subtract 1 from both sides 2x >= 45 -- divide by 2 x >= 22.5 x + 1 = 22.5 + 1 = 23.5 are you sure you wrote this down correctly....because I am getting decimal numbers

OpenStudy (anonymous):

Umm let me check i will repost the problem.

OpenStudy (texaschic101):

does it say consecutive EVEN numbers ?

OpenStudy (texaschic101):

or consecutive ODD numbers ?

OpenStudy (texaschic101):

or just numbers ?

OpenStudy (anonymous):

It dosent say anything.. It says the sum of two consecutive integers is at least 46. What is the least possible pair of integers? and that the answer is 23 & 24

OpenStudy (texaschic101):

oh..I see...they rounded the numbers x >= 22.5 rounds to x >= 23 x + 1 >= 22.5 + 1 = 23.5 rounds to x + 1 >= 24 understand ?

OpenStudy (anonymous):

Ohhhhh okay, got it! Thank you so much. Yeah I was getting the same answer too! They should really say something. Thank you though:)

OpenStudy (texaschic101):

your welcome...your right, they should have mentioned rounding